OpenStudy (anonymous):

I have the answer can someone check me. In the 1940's the population of the tribbles on the island of Kent could be approximated according to the formula P=573(1.021)^t Where t is the number os years since 1940. Find approximately how fast(in tribbles/yeat) the population was growing in 1942.

OpenStudy (anonymous):

Take the derivative with respect to t. \[\frac{ d }{ dt } e^{at} = ae^{at}\] What you can do is express \[1.021 = e^{something}\]

OpenStudy (anonymous):

so would the answer be 1.9326x10^20

OpenStudy (anonymous):

No, you made two mistakes. One is that you used the wrong t; t is the years SINCE 1940.

OpenStudy (anonymous):

so the the t=2

OpenStudy (anonymous):

so then the answer would be 597.319

OpenStudy (anonymous):

No, you haven't take the derivative properly.

OpenStudy (anonymous):

I am confussed on what to do I guess

OpenStudy (anonymous):

\[P = 573e^{at}\] You seem to have correctly found a to be ln(1.021). What is \[\frac{ d }{ dt } 573e^{at}\] ?

OpenStudy (anonymous):

397.173

OpenStudy (anonymous):

What is it as a function of t?

OpenStudy (anonymous):

I am confussed

OpenStudy (anonymous):

o what about 4233.929

OpenStudy (anonymous):

oh no I got it I think is the answer 838.929

OpenStudy (anonymous):

Look at my first post, it gives the rule for finding the derivative.

OpenStudy (anonymous):

ok I think I got it 2.062

OpenStudy (anonymous):

is that answer correct

OpenStudy (anonymous):

can you please tell me if my answer is right!!!!

OpenStudy (anonymous):

It's not what I got. Try to find the expression for the rate of tribble growth\[\frac{ dP }{ dt }\] Then see what the value of dP/dt is when t = 2.